OpenStudy (anonymous):

What are the solutions of x2 + 5x − 24 = 0? Show all work for full credit.

OpenStudy (anonymous):

x² + 5x - 24 = 0 (x + 8)(x - 3) = 0 {factored into two binomials} x + 8 = 0 or x - 3 = 0 {set each factor equal to 0} x = -8 or x = 3 {solved each equation for x}

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hit the link and give me best response